A small vegetable garden has an area of 80 square feet. Its length is 2 feet more than the width. Find the dimensions of the garden
Answer by checkley79(3341)

L=W+2
AREA=L*W
80=(W+2)*W
80=W^2+2W
W^2+2W-80=0
(W+10)(W-8)=0
W-8=0
W=8 ANS. FOR THE WIDTH.
L=8+2=10 ANS. FOR THE LENGTH.
PROOF:
80=10*8
80=80